# The multivalued inverse of a quasiregular map is a quasiregular curve

### v1

@@
Abstract
We observe that a proper quasiregular map $f:\Omega\to\mathbb{R}^{n}$ of finite degree $d$ has a natural multivalued inverse $\accentset{\leftarrow}{f}:f(\Omega)\to\mathcal{A}_{d}(\mathbb{R}^{n})$ taking values in the space $A_{d}(\mathbb{R}^{n})$ of unordered $d$-tuples in $\mathbb{R}^{n}$, and show that the multivalued inverse is a quasiregular curve (in the sense of Pankka) with respect to a natural $n$-form on $\mathcal{A}_{d}(\mathbb{R}^{n})$ (suitably interpreted). We moreover show that $\accentset{\leftarrow}{f}$ is a geometrically quasiconformal map onto its image $\Omega_{f}=\accentset{\leftarrow}{f}(f(\Omega))$, and that the Hausdorff measure on $\Omega_{f}$ is upper Ahlfors regular.
These results demonstrate a new connection between quasiregular maps, quasiregular curves, and quasiconformal theory in metric spaces. To establish them, we develop a theory of pull-backs of differential forms on $\mathcal{A}_{d}(\mathbb{R}^{n})$ by multi-valued Sobolev maps, which may be of independent interest.
1 Introduction
1.1 Overview
@@
$\|Df\|^{n}\leq K\mathbf{J}f\quad\textrm{ almost everywhere},$ (1.2)
@@
where $\mathbf{J}f=\det(Df)$ and $Df$ is the weak differential of $f$. Quasiregular maps have a rich theory [ric93, HKM06], and enjoy many of the same analytic properties as quasiconformal (QC) maps. While they are not necessarily homeomorphisms, a seminal result of Reshetnyak [res67] states that QR maps are open and discrete, and thus in particular locally injective outside a singular branch set of topological dimension at most n2n-2 [vai66]. We call an open discrete map a branched cover.
@@
Quasiregular curves (QR curves) in turn were introduced by Pankka [pankka20], and they generalize quasiregular maps by allowing the dimension of the target to be higher than that of the domain; for example pseudoholomorhic curves [gro85] are QR curves. Much of the classical theory of QR maps extends to QR curves [pankka20, pan-onn21], with one notable exception: Reshetnyak’s theorem. Namely, QR curves need not be branched covers. Indeed, if $k>n$, a QR curve cannot be open, while the more non-trivial failure of discreteness follows from an example in [IVV02].
@@
Compared to QR maps, one of the challenges in QR curves is understanding how the choice of $\omega$ reflects the behaviour of $f$, see for example [hei-pan-pry23, iko-pan24] for some rigidity results and connections with calibrations.
@@
In this paper, we identify a fruther connection between QR maps and QR curves: we construct a multi-valued inverse $\accentset{\leftarrow}{f}$ for a QR map $f:\Omega\to\mathbb{R}^{n}$ of finite degree $d>0$, and show that $\accentset{\leftarrow}{f}$ is naturally a QR curve (see Theorem 1.2). In particular, QR maps have a local (multi-valued) inverse across the singular branch set (and not only outside it) with suitable Sobolev regularity. Earlier, related constructions include a push-forward operator [HKM06, p. 263], [teripekka, Section 4] and a generalized local inverse [ric93, Chapter II.5], [onn-raj09, Section 8], both of which are important for obtaining modulus inequalities.
@@
Our approach is based on Almgren’s space of unordered tuples [del11] (also known as the symmetric product, cf. [hat02, Chapter 4.K]), and brings a new perspective to quasiregular theory; together with a pull-back theory by multi-valued maps developed in this paper (see Theorems 1.1 and 4.7) we obtain e.g. higher integrability properties of the generalized inverse [ric93, Chapter II.5] (see Corollary 1.3 and Theorem 4.15). We describe the construction and briefly discuss some its advantages over alternative approaches below.
1.2 Multi-valued inverse
@@
Given a non-injective (proper) branched cover $f:\Omega\to\mathbb{R}^{n}$, there are several natural ways to interpret the set valued “inverse” map $f(\Omega)\ni y\mapsto f^{-1}(y)$. One is to consider the natural factorization $f=\bar{f}\circ P_{f}$, where $P:\Omega\to\widetilde{\Omega}_{f}:=\Omega/\sim$ is the projection onto the quotient space given by the equivalence relation $x\sim y$ if $f(x)=f(y)$, and $\bar{f}:\widetilde{\Omega}_{f}\to f(\Omega)$ the induced map, which in this case is a homeomorphism and thus invertible. Another approach is to view the set valued inverse as a map $f(\Omega)\to\mathcal{H}(\Omega)$ into the Hausdorff space of non-empty compact subsets of $\Omega$. A drawback of the first approach is that the target space of the inverse depends on $f$ and, more crucially, that $\bar{f}$ need not be geometrically quasiconformal unless $f$ is a BLD-map. The second approach, on the other hand, does not see the multiplicities in the preimage, and $\mathcal{H}(\Omega)$ does not carry any natural differential forms, nor any bi-Lipschitz embedding in Euclidean space (cf. [del11, Theorem 2.1]).
@@
Instead, we give a construction based on Almgren’s space and local degree theory. Let $Q\geq 1$ be a natural number. The Almgren space $\mathcal{A}_{Q}(\mathbb{R}^{n})$ of unordered $Q$-tuples in $\mathbb{R}^{n}$ was used by Almgren in his seminal work on regularity of minimal currents [alm2000], and was further explored by De Lellis et. al. [del11]. It is a $nQ$-dimensional smooth orbifold, but not a manifold (unless $Q=1$), defined as
@@
We refer the reader to Section 2.1 (see in particular Definition 2.1) for the notation and further details used here.
1.3 Statement of results
@@
$\operatorname{tr}(\operatorname{vol}_{\mathbb{R}^{n}}):=\sum_{j}^{d}P_{j}^{*}(\operatorname{vol}_{\mathbb{R}^{n}})$ (1.4)
@@
In particular, if $f:U\to\mathcal{A}_{Q}(\mathbb{R}^{n})$ is a multi-valued (locally) Lipschitz map, the pull-back $f^{*}\omega$ of any $S_{Q}$-invariant smooth $k$-form is a (locally)flat $k$-form on $U$ in the sense of Whitney (cf. Theorem 4.7).
Applying the definition of pull-back in the QR context we can show that, if $f:\Omega\to\mathbb{R}^{n}$ is a quasiregular map, the multi-valued inverse is a quasiregular curve.
@@
Theorems 1.1 and 1.2 together with standard arguments found e.g. in [pan-onn21] yield in particular the higher integrability of the generalized inverse of [ric93, Chapter II.5].
@@
Let $f:\Omega\to\mathbb{R}^{n}$ be a proper quasiregular map of finite degree $d$. Then there exists $p>n$ such that the generalized inverse
$g:f(\Omega)\to\mathbb{R}^{n},\quad g(y):=\sum_{x\in f^{-1}(y)}\iota(f,x)x$
belongs to $W^{1,p}_{loc}(f(\Omega),\mathbb{R}^{n})$.
@@
1.4 Main ideas and further discussion
Theorem 1.1 follows by an approximation argument from it Lipschitz counterpart, Theorem 4.7, whose proof uses induction, decomposition of multi-valued maps (cf. [del11, Proposition 1.6]), and a somewhat delicate approximation scheme. We outline the main idea.
@@ 1.4 Main ideas and further discussion
which is $\frac{1}{\sqrt{d}}$-Lipschitz, see Lemma 4.13 below. The equality $f|_{F}=d\llbracket b(f)\rrbracket|_{F}$ implies that the differentials of the two functions agree almost everywhere. However, since $F$ is closed and not open, we need an approximation argument for the induction step to work. Our approximation scheme consists of a careful interpolation between $f$ and $d\llbracket b(f)\rrbracket$ in a neighbourhood of $F$.
The main step in the proof of Theorem 1.2 are the Sobolev estimates, which we obtain in Theorem 3.7. Our argument here uses Vitali’s covering theorem, which is valid in high generality, and thus Theorem 3.7 probably holds e.g. in the setting of [onn-raj09].
The proof of Theorem 1.4 employs a similar induction argument as in Theorem 4.7 together with some metric quasiconformal theory. We remark here that a large part of metric quasiconformal theory is based on the Ahlfors regularity of the spaces under consideration. Indeed, in our setting this property would imply the Loewner property of the image set $\Omega_{f}:=\accentset{\leftarrow}{f}(f(\Omega)$ by a (deep) result of Semmes [sem96], which in turn implies the equivalence of the various notions of quasiconformality in the metric setting [hei98]. In this paper, we establish the upper Ahlfors $n$-regularity of $\Omega_{f}$ (Proposition 5.2), and an infinitesimal Poincaré inequality. It would be interesting to analyze the geometry of $\Omega_{f}$ further, in particular whether $\Omega_{f}$ admits local bi-Lipschitz parametrizations. By Almgren’s bi-Lipschitz embedding $\mathcal{A}_{d}(\mathbb{R}^{n})\hookrightarrow\mathbb{R}^{N}$, $\Omega_{f}$ can be regarded as a subset of Euclidean space, and the existence of bi-Lipschitz parametrizations is linked with the existence of Cartan–Whitney presentations in a suitable Sobolev class [hei11, [Theorem 1.2].
2 Preliminaries
2.1 Branched covers and quasiregular maps
For the notion of local degree and other topological preliminaries we refer the reader to [ric93, Chapter I.4]. In this paper we call a continuous open discrete map a branched cover. A precompact open set $U\subset X$ is a normal neighbourhood of $f$ if $\partial f(U)=f(\partial U)$. A branched cover $f:X\to Y$ between topological $n$-manifolds has a well-defined local index $\iota(f,x)$ for every $x\in X$, satisfying
@@ 2.1 Branched covers and quasiregular maps
is a closed set of topological dimension at most n2n-2 [vai66]. It follows that branched covers are either sense preserving ($\iota(f,\cdot)\geq 1$ everywhere) or sense reversing ($\iota(f,\cdot)\leq-1$ everywhere). If $f:X\to Y$ is a proper branched cover of degree finite $d>0$, then
@@
Recall that a Sobolev map $f\in W^{1,n}_{loc}(\Omega,\mathbb{R}^{n})$ ($\Omega\subset\mathbb{R}^{n}$ open, $n\geq 2$) is called quasiregular, if there exists $K$ such that (1.2) holds. See [ric93] for an excellent introduction to quasiregular maps. The least constant $K$ in (1.2) is called the outer distortion and denoted $K_{O}(f)$. The inner distortion $K_{I}(f)$ is the smallest constant $K$ such that
$\mathbf{J}f\leq K\min_{v\in S^{n-1}}\{\|Df(v)\|^{n}\}\ \textrm{ almost everywhere}.$
@@
By the seminal result of Reshetnyak [res67], non-constant quasiregular maps are continuous, open and discrete. In particular they are sense-preserving and $\mathbf{J}f\in(0,\infty)$ almost everywhere.
2.2 Space of unordered tuples
@@
$d_{\mathcal{A}}(\bar{x},\bar{y})=\min_{\sigma\in\Sigma_{d}}\Big(\sum_{j}^{d}d(x_{j},y_{\sigma(j)})^{2}\Big)^{1/2},$ (2.2)
for $\bar{x}=\llbracket x_{1},\ldots,x_{d}\rrbracket,\ \bar{y}=\llbracket y_{1},\ldots,y_{d}\rrbracket$. Here, the action by isometries on $X^{d}$ of the symmetric group $S_{d}$ is given by $\sigma\cdot(x_{1},\ldots,x_{d})=(x_{\sigma^{-1}(1)},\ldots,x_{\sigma^{-1}(d)})$. The space $\mathcal{A}_{d}(X)$ can alternatively be described as a subset of the Wasserstein space $\mathcal{P}_{2}(X)$, consisting of probability measures $\mu\in\mathcal{P}_{2}(X)$ such that $\operatorname{spt}(\mu)\leq d$ (note however that the metric inherited from $\mathcal{P}_{2}(X)$ differs from (2.2) by a factor of $1/\sqrt{d}$). Following [del11], we denote elements of $\mathcal{A}_{d}(X)$ either by $\llbracket x_{1},\ldots,x_{d}\rrbracket$ of by $\sum_{j}^{m}a_{j}\llbracket x_{j}\rrbracket$, where the numbers $a_{j}>0$ satisfy $\sum_{j}^{m}a_{j}=d$.
The following Rademacher’s theorem for multi-valued Lipschitz maps is from [del11, Theorem 1.13]. In the statement $m,n,d\geq 1$ are natural numbers and $U\subset\mathbb{R}^{m}$ an open set.
@@
Let $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ be locally Lipschitz. Then $f$ is differentiable for almost every $x_{0}\in U$ in the following sense: there are linear maps $L_{1},\ldots,L_{d}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ so that
@@
The differential $Df:=\llbracket L_{1},\ldots,L_{d}\rrbracket$ can be regarded as an element in $\mathcal{A}_{d}(\mathbb{R}^{n\times n})$, but note that this does not determine it unambiguously, cf. [del11, Remark 1.11]. If $f_{i}:U\to\mathcal{A}_{d_{i}}(\mathbb{R}^{n})$, $i=0,1$, are differentiable at $x\in U$ and $d_{0}+d_{1}=d$, then the map $f:=\llbracket f_{0},f_{1}\rrbracket:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is differentiable at $x\in U$, and $D_{x}f=\llbracket D_{x}f_{0},D_{x}f_{1}\rrbracket$.
We refer the reader to [del11, Chapter 2.2 and Chapter 4] for the theory of Sobolev maps $W^{1,p}_{loc}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$ and mention here that Sobolev maps are approximately differentiable almost everywhere [del11, Corollary 2.7]. This follows from the following Lusin-type approximations by Lipschitz maps [del11, Proposition 2.5].
@@
The claim is stated with $\{f\neq f_{\lambda}\}$ replaced by $U$ on the right hand side of (2.3) in [del11, Proposition 2.5], but the proof and references therein yield the inequality stated here. ∎
3 Multivalued inverse
3.1 Basic properties for branched covers
Throughout this subsection, $f:X\to Y$ is a branched cover of degree $d$ between metric spaces $X,Y$ homeomorphic to topological $n$-manifolds, $n\geq 2$. We start with a simple lemma.

### v2

# Pull-back of differential forms by multi-valued Sobolev maps, and the quasiregularity of the multi-valued inverse of a quasiregular map

@@
Abstract.
We use Almgren’s framework of multi-valued maps to construct a multi-valued inverse $\accentset{\leftarrow}{f}:f(\Omega)\to\mathcal{A}_{d}(\mathbb{R}^{n})$ of a quasiregular map $f:\Omega\to\mathbb{R}^{n}$ of finite degree $d$. We then develop a pull-back theory of differential forms on $\mathcal{A}_{d}(\mathbb{R}^{n})$ by Sobolev maps, and use it to show that the multi-valued inverse is a quasiregular $\omega$-curve (in the sense of Pankka) with respect to a natural $n$-form $\omega$ (suitably interpreted). The pull-back theory is of independent interest, and allows us to conclude e.g. higher Sobolev integrability and quasiminimality of the multi-valued inverse.
Key words and phrases:
Quasiregular maps, quasiregular curves, multi-valued maps, Almgren space, Sobolev differential forms, weak exterior derivative
2020 Mathematics Subject Classification:
30C65, 35R70, 53C23, 30L10
1. Introduction
1.1. Overview
@@
$\|Df\|^{n}\leq KJ_{f}\quad\textrm{ almost everywhere},$ (1.2)
@@
where $J_{f}=\det(Df)$ and $Df$ is the weak differential of $f$. Quasiregular maps have a rich theory [26, 9], and enjoy many of the same analytic properties as quasiconformal (QC) maps. While they are not necessarily homeomorphisms, a seminal result of Reshetnyak [25] states that QR maps are open and discrete, and thus in particular locally injective outside a singular branch set of topological dimension at most n2n-2 [28]. We call an open discrete map a branched cover.
@@
Quasiregular curves (QR curves) in turn were introduced by Pankka [24], and they generalize quasiregular maps by allowing the dimension of the target to be higher than that of the domain; for example pseudoholomorhic curves [5] are QR curves. Much of the classical theory of QR maps extends to QR curves [24, 21], with one notable exception: Reshetnyak’s theorem. Namely, QR curves need not be branched covers. Indeed, if $k>n$, a QR curve cannot be open, while the more non-trivial failure of discreteness follows from an example in [16].
@@
Compared to QR maps, one of the challenges in QR curves is understanding how the choice of $\omega$ reflects the behaviour of $f$, see for example [7, 13] for some rigidity results and connections with calibrations.
@@
In this paper, we identify a fruther connection between QR maps and QR curves: we construct a multi-valued inverse $\accentset{\leftarrow}{f}$ for a QR map $f:\Omega\to\mathbb{R}^{n}$ of finite degree $d>0$, and show that $\accentset{\leftarrow}{f}$ is naturally a QR curve (see Theorem 1.2). In particular, QR maps have a local (multi-valued) inverse across the singular branch set (and not only outside it) with suitable Sobolev regularity. Earlier, related constructions include a push-forward operator [9, p. 263], [23, Section 4] and a generalized local inverse [26, Chapter II.5], [22, Section 8], both of which are important for obtaining modulus inequalities.
@@
Our approach is based on Almgren’s space of unordered tuples [2] (also known as the symmetric product, cf. [6, Chapter 4.K]), and brings a new perspective to quasiregular theory; together with a pull-back theory by multi-valued maps developed in this paper (see Theorems 1.1 and 4.7) we obtain e.g. higher integrability of the generalized inverse [26, Chapter II.5] (see Corollary 1.3 and Theorem 5.3) as well as quasiminimality of the multi-valued inverse (see Theorem 5.4). We describe the construction and briefly discuss some its advantages over alternative approaches below.
1.2. Multi-valued inverse
@@
Given a non-injective (proper) branched cover $f:\Omega\to\mathbb{R}^{n}$, there are several natural ways to interpret the set valued “inverse” map $f(\Omega)\ni y\mapsto f^{-1}(y)$. One is to consider the natural factorization $f=\bar{f}\circ P_{f}$, where $P_{f}:\Omega\to\widetilde{\Omega}_{f}:=\Omega/\sim$ is the projection onto the quotient space given by the equivalence relation $x\sim y$ if $f(x)=f(y)$, and $\bar{f}:\widetilde{\Omega}_{f}\to f(\Omega)$ the induced map, which in this case is a homeomorphism and thus invertible. Another approach is to view the set valued inverse as a map $f(\Omega)\to\mathcal{H}(\Omega)$ into the Hausdorff space of non-empty compact subsets of $\Omega$. A drawback of the first approach is that the target space of the inverse depends on $f$ and, more crucially, that $\bar{f}$ need not be geometrically quasiconformal unless $f$ is a BLD-map. The second approach, on the other hand, does not detect the multiplicities in the preimage, and $\mathcal{H}(\Omega)$ does not carry any natural differential forms, nor any bi-Lipschitz embedding in Euclidean space (cf. [2, Theorem 2.1]).
@@
Instead, we give a construction based on Almgren’s space and local degree theory. Let $Q\geq 1$ be a natural number. The Almgren space $\mathcal{A}_{Q}(\mathbb{R}^{n})$ of unordered $Q$-tuples in $\mathbb{R}^{n}$ was used by Almgren in his seminal work on regularity of minimal currents [1], and was further explored by De Lellis et. al. [2]. It is a $nQ$-dimensional smooth orbifold, but not a manifold (unless $Q=1$), defined as
@@
We refer the reader to Section 2.1 (see in particular Definition 2.1) for the notation and further details.
1.3. Statement of results
@@
$\operatorname{tr}(\operatorname{vol}_{\mathbb{R}^{n}}):=\sum_{j}^{Q}P_{j}^{*}(\operatorname{vol}_{\mathbb{R}^{n}})$ (1.4)
@@
In particular, if $f:U\to\mathcal{A}_{Q}(\mathbb{R}^{n})$ is a multi-valued (locally) Lipschitz map, the pull-back $f^{*}\omega$ of any $S_{Q}$-invariant smooth $k$-form $\omega$ is a (locally) flat $k$-form on $U$ in the sense of Whitney (cf. Theorem 4.7).
Applying the definition of pull-back in the QR context we show that, if $f:\Omega\to\mathbb{R}^{n}$ is a quasiregular map, the multi-valued inverse is a quasiregular curve.
@@
Here $K_{I}(f)$ is the inner distortion of the quasiregular map $f$, see Section 2.1. Theorems 1.1 and 1.2 together with standard arguments found e.g. in [21] yield in particular the higher integrability of the generalized inverse of [26, Chapter II.5]. Recall that, given a proper quasiregular map $f:\Omega\to\mathbb{R}^{n}$, the generalized inverse is the map $g_{f}:f(\Omega)\to\mathbb{R}^{n}$ given by
$g:f(\Omega)\to\mathbb{R}^{n},\quad g(y):=\sum_{x\in f^{-1}(y)}\iota(f,x)x.$ (1.5)
We record this higher integrability result in Corollary 1.3 below, along with the somewhat surprising fact that the generalized inverse (1.5) is also a quasiminimizer of the Dirichlet energy. While both statements follow from the corresponding results for the multi-valued inverse (Theorems 5.3 and 5.4), the latter requires further analysis of the structure of Almgren’s space, presented in Appendix A. Indeed in Appendix A we show that $\mathcal{A}_{d}(\mathbb{R}^{n})$ splits off a factor $\mathbb{R}^{n}$, and cannot split off any higher dimensional Euclidean factors, cf. Proposition A.1.
@@
Let $f:\Omega\to\mathbb{R}^{n}$ be a proper quasiregular map, and $g_{f}$ its generalized inverse. Then
  • (i)
    there exists $p>n$ such that $g_{f}\in W^{1,p}_{loc}(f(\Omega),\mathbb{R}^{n})$;
  • (ii)
    if $h\in W^{1,n}_{loc}(f(\Omega),\mathbb{R}^{n})$ is such that $\{h\neq g_{f}\}$ is essentially contained in a compact set $K\subset f(\Omega)$, then
    $\int_{K}\|Dg_{f}\|^{n}\mathrm{d}x\leq d^{n/2-1}K_{I}\int_{K}\|Dh\|^{n}\mathrm{d}x.$
@@
1.4. Main ideas and further discussion
Theorem 1.1 follows by an approximation argument from its Lipschitz counterpart, Theorem 4.7, whose proof uses induction, decomposition of multi-valued maps (cf. [2, Proposition 1.6]), and a somewhat delicate approximation scheme. We outline the main idea.
@@ 1.4. Main ideas and further discussion
which is $\frac{1}{\sqrt{d}}$-Lipschitz, see Corollary A.2. The equality $f|_{F}=d\llbracket b(f)\rrbracket|_{F}$ implies that the differentials of the two functions agree almost everywhere on $F$. However, since $F$ is closed and not open, we need an approximation argument for the induction step to work. Our approximation scheme consists of a careful interpolation between $f$ and $d\llbracket b(f)\rrbracket$ in a neighbourhood of $F$.
The main step in the proof of Theorem 1.2 are the Sobolev estimates, which we obtain in Theorem 3.7. Our argument here uses Vitali’s covering theorem, which is valid in high generality, and thus Theorem 3.7 probably holds e.g. in the setting of [22].
The proof of Theorem 1.4 employs a similar induction argument as in Theorem 4.7 together with some metric quasiconformal theory. We remark here that a large part of metric quasiconformal theory is based on the Ahlfors regularity of the spaces under consideration. Indeed, in our setting this property would imply the Loewner property of the image set $\Omega_{f}:=\accentset{\leftarrow}{f}(f(\Omega)$ by a (deep) result of Semmes [27], which in turn implies the equivalence of the various notions of quasiconformality in the metric setting [11]. In this paper, we establish the upper Ahlfors $n$-regularity of $\Omega_{f}$ (Proposition 6.2), and an infinitesimal Poincaré inequality. It would be interesting to analyze the geometry of $\Omega_{f}$ further, in particular whether $\Omega_{f}$ admits local bi-Lipschitz parametrizations. By Almgren’s bi-Lipschitz embedding $\mathcal{A}_{d}(\mathbb{R}^{n})\hookrightarrow\mathbb{R}^{N}$, $\Omega_{f}$ can be regarded as a subset of Euclidean space, and the existence of bi-Lipschitz parametrizations is linked with the existence of Cartan–Whitney presentations in a suitable Sobolev class [8, [Theorem 1.2].
Acknowledgements
The research in this manuscript was supported by Research Council of Finland grant no. 355122. The author thanks Pekka Pankka, Ilmari Kangasniemi and Toni Ikonen for useful discussions on the manuscript.
2. Preliminaries
2.1. Branched covers and quasiregular maps
For the notion of local degree and other topological preliminaries we refer the reader to [26, Chapter I.4]. In this paper we call a continuous open discrete map a branched cover. A precompact open set $U\subset X$ is a normal neighbourhood of $f$ if $\partial f(U)=f(\partial U)$. A branched cover $f:X\to Y$ between topological $n$-manifolds has a well-defined local index $\iota(f,x)$ for every $x\in X$, satisfying
@@ 2.1. Branched covers and quasiregular maps
is a closed set of topological dimension at most n2n-2 [28]. It follows that branched covers are either sense preserving ($\iota(f,\cdot)\geq 1$ everywhere) or sense reversing ($\iota(f,\cdot)\leq-1$ everywhere). If $f:X\to Y$ is a proper branched cover of degree finite $d>0$, then
@@
Recall that a Sobolev map $f\in W^{1,n}_{loc}(\Omega,\mathbb{R}^{n})$ ($\Omega\subset\mathbb{R}^{n}$ open, $n\geq 2$) is called quasiregular, if there exists $K$ such that (1.2) holds. See [26] for an excellent introduction to quasiregular maps. The least constant $K$ in (1.2) is called the outer distortion and denoted $K_{O}(f)$. The inner distortion $K_{I}(f)$ is the smallest constant $K$ such that
$J_{f}\leq K\min_{v\in S^{n-1}}\{\|Df(v)\|^{n}\}\ \textrm{ almost everywhere}.$
@@
By the seminal result of Reshetnyak [25], non-constant quasiregular maps are continuous, open and discrete. In particular they are sense-preserving and $J_{f}\in(0,\infty)$ almost everywhere.
2.2. Space of unordered tuples
@@
$d_{\mathcal{A}}(\bar{x},\bar{y})=\min_{\sigma\in S_{d}}\Big(\sum_{j}^{d}d(x_{j},y_{\sigma(j)})^{2}\Big)^{1/2},$ (2.2)
for $\bar{x}=\llbracket x_{1},\ldots,x_{d}\rrbracket,\ \bar{y}=\llbracket y_{1},\ldots,y_{d}\rrbracket$. Here, the action by isometries on $X^{d}$ of the symmetric group $S_{d}$ is given by $\sigma\cdot(x_{1},\ldots,x_{d})=(x_{\sigma^{-1}(1)},\ldots,x_{\sigma^{-1}(d)})$. The space $\mathcal{A}_{d}(X)$ can alternatively be described as a subset of the Wasserstein space $\mathcal{P}_{2}(X)$, consisting of probability measures $\mu\in\mathcal{P}_{2}(X)$ such that $\operatorname{spt}(\mu)\leq d$ and $d\cdot\mu(\{x\})\in\mathbb{N}$ for all $x\in X$ (note however that the metric inherited from $\mathcal{P}_{2}(X)$ differs from (2.2) by a factor of $1/\sqrt{d}$). Following [2], we denote elements of $\mathcal{A}_{d}(X)$ either by $\llbracket x_{1},\ldots,x_{d}\rrbracket$ or by $\sum_{j}^{m}a_{j}\llbracket x_{j}\rrbracket$, where the numbers $a_{j}>0$ satisfy $\sum_{j}^{m}a_{j}=d$.
Multi-valued differentiability
The following Rademacher’s theorem for multi-valued Lipschitz maps is from [2, Theorem 1.13]. In the statement $m,n,d\geq 1$ are natural numbers and $U\subset\mathbb{R}^{m}$ an open set.
@@
Let $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ be locally Lipschitz. Then $f$ is differentiable for almost every $x_{0}\in U$ in the following sense: there are linear maps $L_{1},\ldots,L_{d}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ so that
@@
The differential $Df:=\llbracket L_{1},\ldots,L_{d}\rrbracket$ can be regarded as an element in $\mathcal{A}_{d}(\mathbb{R}^{n\times n})$, but note that this does not determine it unambiguously, cf. [2, Remark 1.11]. If $f_{i}:U\to\mathcal{A}_{d_{i}}(\mathbb{R}^{n})$, $i=0,1$, are differentiable at $x\in U$ and $d_{0}+d_{1}=d$, then the map $f:=\llbracket f_{0},f_{1}\rrbracket:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is differentiable at $x\in U$, and $D_{x}f=\llbracket D_{x}f_{0},D_{x}f_{1}\rrbracket$.
Multi-valued Sobolev maps
We refer the reader to [2, Chapter 2.2 and Chapter 4] for the theory of Sobolev maps $W^{1,p}_{loc}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$ and mention here that Sobolev maps are approximately differentiable almost everywhere [2, Corollary 2.7]. This follows from the following Lusin-type approximations by Lipschitz maps [2, Proposition 2.5].
@@
The claim is stated with $\{f\neq f_{\lambda}\}$ replaced by $U$ on the right hand side of (2.3) in [2, Proposition 2.5], but the proof and references therein yield the inequality stated here. ∎
Metric Jacobian
Let $s$ be a seminorm on $\mathbb{R}^{n}$. The Jacobian $\mathbf{J}(s)$ is the unique number satisfying $\mathcal{H}^{n}_{s}=\mathbf{J}(s)\mathcal{L}^{n}$, where $\mathcal{H}^{n}_{s}$ is zero if $s$ is degenerate, and the Hausdorff $n$-measure with respect to the metric induced by $s$ otherwise. If $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is (approximately) differentiable at $x\in U$, we denote
$\mathbf{J}f(x)=\mathbf{J}(\operatorname{md}_{x}f),$
where $\operatorname{md}_{x}f$ is the seminorm on $\mathbb{R}^{n}$ given by $\operatorname{md}_{x}f(v)=|D_{x}f(v)|$ for $v\in\mathbb{R}^{n}$. [^1] Note that this is well-defined since the norm of $D_{x}f(v)$ is independent of the permutation of order of the elements in $D_{x}f=(D_{x}f_{1},\ldots,D_{x}f_{d})$. We refer to [18] where the metric differential and Jacobian were introduced in order to prove an area formula for maps into metric spaces (see also [17]). We record the following identity for future use.
Lemma 2.4.
Suppose $L=\llbracket L_{1},\ldots,L_{d}\rrbracket:\mathbb{R}^{n}\to\mathcal{A}_{d}(\mathbb{R}^{n})$, where $L_{1},\ldots,L_{d}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ are linear maps, and let $s$ be the seminorm given by $s(v)=\Big(\sum_{j}^{d}|L_{j}(v)|^{2}\Big)^{1/2}$ for $v\in\mathbb{R}^{n}$. Then
$\mathbf{J}(s)=\det(L_{1}^{T}L_{1}+\cdots+L_{d}^{T}L_{d})^{1/2}.$
In particular, if $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is differentiable a.e., then
$\mathbf{J}f=\det\Big(\sum_{j}^{d}Df_{j}^{T}Df_{j}\Big).$
Proof.
If $f$ is differentiable almost everywhere, then it is metrically differentiable almost everywhere and
$\operatorname{\operatorname{md}}_{x}f(v)=\Big(\sum_{j}^{d}|Df_{j}(v)|^{2}\Big)^{1/2},\quad v\in\mathbb{R}^{m}$
for a.e. $x\in U$. Thus the second claim readily follows from the first.
To prove the first, it suffices to consider the linear map $L=(L_{1},\ldots,L_{d}):\mathbb{R}^{n}\to(\mathbb{R}^{n})^{d}$. By the area formula on Euclidean spaces, we have that $J_{n}(L)L_{\ast}\mathcal{L}^{n}=\mathcal{H}^{n}|_{\operatorname{Im}(L)}$, where
$J_{n}(L)=|\det(L^{T}L)|^{1/2}=|\det(L_{1}^{T}L_{1}+\cdots+L_{d}^{T}L_{d})|^{1/2}$
Now $J_{n}(L)\neq 0$ if and only if $L$ has full rank, in which case $s$ is a genuine norm. Note that $L:(\mathbb{R}^{n},s)\to(\operatorname{Im}(L),\|\cdot\|_{Eucl})$ is an isometry, and thus $L_{\ast}\mathcal{H}^{n}_{s}=\mathcal{H}^{n}|_{\operatorname{Im}(L)}$. Since $L:\mathbb{R}^{n}\to\operatorname{Im}(L)$ is bijective, this yields $\mathcal{H}^{n}_{s}=J_{n}(L)\mathcal{L}^{n}$. ∎
3. Multivalued inverse
3.1. Basic properties for branched covers
Throughout this subsection, $f:X\to Y$ is a proper branched cover of degree $d$ between metric spaces $X,Y$ homeomorphic to topological $n$-manifolds, $n\geq 2$. We start with a simple lemma.